# Trust-Preserved Human-Robot Shared Autonomy enabled by Bayesian Relational Event Modeling

### v2

@@ IV-B Coordination between Trust and Autonomy Levels
In real-world human-robot teaming scenarios, the collaboration between humans and robots is a long-term dynamic process, where all team members continuously engage in the process of establishing and calibrating trust among each other. Therefore, it is expected for robots to deploy a trust-preserve strategy to actively establish, maintain, and repair human trust, rather than merely passively adapting to it.
IV-C Trust-preserved Shared Autonomy
Methods for trust repair
Description
Apology
Admission of fault and statement of remorse
Denial
Recognition of error and assertion that they are not at fault
Control
Statement of control/awareness of situation
Convey uncertainty directly
Conveying uncertainty directly by providing machine confidence indicators
Show critical states
Enhance an operator’s mental state by showing only the critical states, situations that require immediate action, of an algorithm’s policy
@@ IV-C Trust-preserved Shared Autonomy
To devise a trust-preserved shared autonomy strategy, we resort to an analysis of the temporal dynamics about trust formation and maintenance within a broader perspective of longitudinal teaming. Punctuated Equilibrium Model (PEM) [41], which describes the pattern of evolution in evolutionary biology and paleontology, reveals that the dynamics of groups often involve long periods of relative stability (equilibrium) punctuated by relatively short bursts of rapid change (punctuation), as illustrated in Fig. 6. According to the paradigm of PEM, the team structure is considered stable and experiences limited change over the evolutionary periods, representing equilibrium. These periods are interrupted by brief but significant episodes of revolutionary periods, where the team structure may disassemble, reconfigure, and experience qualitative, metamorphic change. The transition from an evolutionary period into a revolutionary period can be characterized by the inflection point of the team dynamics, which is a critical moment or juncture when a significant change or turning point occurs in the team structure.
In the context of human-robot teaming, evolutionary periods represent the periods when humans and robots have established and maintained a relatively stable trust and relationship. Within those periods, robots can simply adjust the autonomy level to match the estimated trust level $L_{\alpha}\propto L_{\beta}$. However, for revolutionary periods, which often happen at the beginning of the teaming process or when trust violation occurs, additional trust repair approaches should be deployed to establish or re-establish the trust between team members.
There are various kinds of methods to repair the trust in human-robot teams, for example, by explaining the cause and/or situational nature of a mistake, or making promises about future behavior. Following [24], we summarize several typical strategies that are suitable for remote human-robot collaboration scenarios, as demonstrated in Table. I. For instance, in the scenario of collaborative SAR, the initial trust from the human may be low due to unfamiliarity. Then during this initial revolutionary period, the robots could explain their decisions to the human and seek confirmation from the human whenever making a counter-command decision. It is also possible that the robots occasionally make decisions that appear “unethical” to the human, resulting in a loss of trust. In such cases, rather than simply switching to a lower autonomy level, the robots could state their awareness of the situation and assure the human by providing the estimated probability of a dangerous condition.
V User Study
@@
V-A Study Design
Participants: We recruited 16 college student subjects (3 female, average ages 26.06$\pm$4.93 years) to participate in our study. All subjects were provided informed consent prior to the experiment.
Tasks: Each participant needs to play a simulated human-robot collaborative SAR game, as described in Section II.
@@ V-A Study Design
The participant first played the game in the practice phase to get familiar with the task and interface. In this phase, the robots were assigned two fixed autonomy levels, one with a higher autonomy level ($L_{\alpha}=0.9$) and one with a lower autonomy level ($L_{\alpha}=0.1$). To prevent the participants from learning how to go about the entire SAR operation, the scenarios (which building was burning, which one had a gas leak, where they were located, which one had people) were randomized for each trial. Participants could try as many trials as they wish at this stage, after which they were asked to report their 10-point Likert scale on their trust in each robot. The interaction and trust data collected at this stage was utilized to perform the offline learning for team properties grounding, which captured the individual differences of the participants (e.g. if a higher trust predicts the probability of sending fewer inquiring messages).
@@ V-B Measures
Subjective Measures: We administered a 10-point Likert scale survey for each condition. The test distinguishes five dimensions of collaboration fluency: How easy it was to complete the tasks; How helpful the robots were to complete the tasks; How much trust had been put into the robots about their situation awareness; In what degree the robots acted like a member of the team; In what degree the robots’ responses helped improve the trust. We also asked each participant’s preference on which condition they would choose to work with in the future.
V-C Results
@@
The objective and subjective data are analyzed, whose means and standard deviations are displayed in Fig. 8 and Fig. 9, respectively.
Figure 8: Means and standard deviations of objective measures. Significant enhancement of success rate and reduction of workload are demonstrated.
Figure 9: Means and standard deviations of subjective measures. There are significant improvements in the users’ experience over all five dimensions of collaboration fluency.
[…]
The reason that there is no significant difference in the task duration between the two conditions may be due to the fact that only the completion time of successful task execution was recorded. In those trials, users may occasionally deploy a near-optimal strategy of task execution, which leads to a near-optimal task completion time.
[…]
VI Conclusion and Future Work
@@
In this paper, a trust-preserved shared autonomy strategy was proposed, which allows robots to seamlessly adjust their autonomy level, striving to optimize team performance and enhance their acceptance among human collaborators. Providing an innovative approach to model the dynamics of human trust based on Bayesian REM, this paper opens a new perspective of trust modeling in remote human-robot interaction scenarios. Future work will investigate the application of the proposed method to more general and complex human-robot teaming scenarios, for example, considering the disuse of robots or dynamically changing priorities.

### v3

@@ VI-B Coordination between Trust and Autonomy
VI-C Trust-preserved Shared Autonomy
In real-world human-robot teaming scenarios, the collaboration between humans and robots is a long-term dynamic process, where all team members continuously engage in the process of establishing and calibrating trust among each other. Therefore, robots are expected to deploy a trust-preserve strategy to actively establish, maintain, and repair human trust, rather than merely passively adapting to it.
On that account, we take a longitudinal perspective of human-robot teaming and analyze the temporal dynamics of trust formation and maintenance. Punctuated Equilibrium Model (PEM) [Sabherwal2001TheModel], which describes the pattern of evolution in evolutionary biology and paleontology, reveals that the dynamics of groups often involve long periods of relative stability (equilibrium) punctuated by relatively short bursts of rapid change (punctuation), as illustrated in Fig. 6. According to the paradigm of PEM, the team structure is considered stable and experiences limited change over the evolutionary periods, representing equilibrium. These periods are interrupted by brief but significant episodes of revolutionary periods, where the team structure may disassemble, reconfigure, and experience qualitative, metamorphic change. In the context of human-robot teaming, evolutionary periods represent the periods when humans and robots have established and maintained a relatively stable trust and relationship. Within those periods, robots can simply adjust the autonomy level to match the estimated trust level. However, for revolutionary periods, which often happen at the beginning of the teaming process or when trust violation occurs, additional trust repair approaches should be deployed to establish or re-establish the trust between team members.
Methods for trust repair Description
Apology Admission of fault and statement of remorse
Denial Recognition of error and assertion that they are not at fault
Control Statement of control/awareness of situation
Convey uncertainty directly Conveying uncertainty directly by providing machine confidence indicators
Show critical states Enhance an operator’s mental state by showing only the critical states, situations that require immediate action, of an algorithm’s policy
@@ VI-C Trust-preserved Shared Autonomy
The transition from an evolutionary period into a revolutionary period is characterized by the inflection point of the team dynamics, which can be detected by inspecting the derivatives of the estimated trust level. Whenever a revolutionary period is detected, various trust repair approaches, including explaining the cause of a mistake, or making promises about future behavior, can be actively deployed. Following [DeVisser2020TowardsTeams], we summarize several strategies suitable for human-robot collaboration SAR scenarios, as demonstrated in Table I. For instance, robots could explain their decisions to the human (e.g., “My priority has been adjusted due to an observed dangerous condition in Building X.”), or seek confirmation from human whenever making a counter-command decision (e.g., “Instruction received. However, I have the following additional environmental information that may be useful for you to refine the decisions. Please resend your decision.”).
VII User Study
@@
VII-A Study Design
[…]
Tasks: Each participant needs to play a simulated human-robot collaborative SAR game, as described in Section IV.
@@ VII-A Study Design
The participant first played the game in the practice phase to get familiar with the task and interface. In this phase, the robots were assigned two fixed autonomy levels, one with a higher autonomy level ($L_{\text{autonomy}}=0.9$) and one with a lower autonomy level ($L_{\text{autonomy}}=0.1$). To prevent the participants from learning how to go about the entire SAR operation, the scenarios (which building was burning, which one had a gas leak, where they were located, which one had people) were randomized for each trial. Participants could try as many trials as they wish at this stage, after which they were asked to report their 11-point Likert scale on their trust in each robot (in what degree they tend to accept the robot’s decisions). The interaction and trust data collected at this stage was utilized to perform the offline learning for team properties grounding, which captured the individual differences of the participants (e.g. if a higher trust predicts the probability of sending fewer inquiring messages).
@@ VII-B Measures
Subjective Measures: We administered an 11-point Likert scale (from “0-Not at all agree” to “10-Totally agree”) survey for each condition, which is a tailored version of [Hoffman2019EvaluatingCollaboration] within the SAR scenario. The test distinguishes five dimensions of collaboration fluency: How easy it was to complete the tasks; How helpful the robots were to complete the tasks; How much trust had been put into the robots about their situation awareness; In what degree the robots acted like a member of the team; In what degree the robots’ responses helped improve the trust. We also asked each participant’s preference on which condition they prefer to work with in the future.
VII-C Results
@@
The means, standard deviations, and t-test results (where n.s., $\ast$, $\ast\ast$, $\ast\ast\ast$, $\ast\ast\ast$ denote $p>0.05,\leq 0.05,\leq 0.01,\leq 0.001,\leq 0.0001$, respectively) of objective and subjective measures are displayed in Fig. 8 and Fig. 9.
Figure 8: Results of objective measures.
Figure 9: Results of subjective measures.
Objective Results: As illustrated in Fig. 8, the proposed Trust-preserved SA approach significantly increased the success rate compared to the Baseline SA approach, $t(30)=6.41,p<.00001$, achieving 98% average success rate. The effect size was medium, with a Cohen’s d of 0.7. In addition, it can also reduce the workload of humans by effectively decreasing the number of commands sent from the human teammate, t(30)=2.62,p=.009t(30)=-2.62,p=.009. The effect size was very small, with a Cohen’s d of 0.05.
Subjective Results: From Fig. 9, it is obvious to interpret significant improvements in the users’ experience over all five dimensions of collaboration fluency. Furthermore, when asked to provide their preferences of conditions to work with in the future, all 16 participants chose our proposed Trust-preserved SA method.
VII-D Discussion
Those results were consistent with our hypothesis that, the proposed trust-preserved shared autonomy strategy can improve the performance of human-robot teaming in terms of both objective measures about task completion, and subjective measures about the human acceptability.
During this study, we observed that humans tended to place low levels of initial trust in robots at the beginning of the task. And the reason that our proposed strategy out-performed the baseline may due to its ability to detect human trust and actively deploy trust repair approaches when necessary. This enabled robots to continuously operate with a high level of autonomy to better facilitate the overall team performance. The reason that there is no significant difference in the task duration between the two conditions may be because only the completion time of successful task execution was recorded. Due to the existence of time limit about gas explosion, task duration larger than certain threshold was recognized as failure and will not be recorded.
Although primitive, the proposed approach and designed user study within the SAR scenario provide a valuable first step to develop trust inference and trust-preserved shared autonomy for remote human-robot teaming. After the user study, we asked the participants to provide their suggestions to enhance the scenario setup. Although 14 of them were satisfied with the design, 2 of them suggested to provide an option of human manual override or disuse, which could ensure human authority when needed. Therefore, one interesting next step is to extend this basic scenario setup to more general and complex human-robot teaming scenarios, for example, considering the disuse of robots or dynamically changing priorities.
VII-E Limitations
The 16 participants in our user study were college students and had no prior experiences on real-world SAR. By contrast, experts with professional training on SAR may have biased expectations and assessments of the robot performance. In addition, the real SAR scenario is far more complicated than the one emulated here, which may lead to different results.
VIII Conclusion and Future Work
@@
In this paper, a trust-preserved shared autonomy strategy was proposed, which allows robots to seamlessly adjust their autonomy level, striving to optimize team performance and enhance their acceptance among human collaborators. Providing an innovative approach to model the dynamics of human trust based on Bayesian REM, this paper opens a new perspective of trust modeling in remote human-robot interaction scenarios. Future work plans to validate the proposed method using experiments performed with real robots and realistic interfaces.